a plane travels from NY to CA at 550mph. A second plane travels 30 minutes later at 625mph. how many hours later will plane two pass the first?

Answer by richwmiller(17219)

Gaining speed 625 -550= 75 mph
Headstart 550*0.5 =275 miles
Time to catch up 275/75=3.66666667 hours
Algebraic solution
550(t+0.5)=625t
550t+275.0=625t
275.0=625t-550t
275.0=75t
t=3.66666667 hours
code drt1